# Counting Hallucinations in Diffusion Models

### v1

@@ A.1 Poof of the Bound of Total Accumulated Distributional Error in Ancestral Sampling
Then, according to the Data Processing Inequality (Cover & Thomas, 2012), marginalizing (or discarding some random variables) will not increase the KL divergence:
@@
where equality (=) would strictly hold only in the ideal case where $q({\bm{x}}_{T})=p_{\theta}({\bm{x}}_{T})$ and $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})=q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})$ for all $t$ and ${\bm{x}}_{t}$, which means the diffusion model can accurately recover the underlying data distribution $q({\bm{x}}_{0})$. This upper bound for the overall KL divergence (${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{0})\|p_{\theta}({\bm{x}}_{0}))$) between true data and model-generated data is identical to the variational bound derived in DDPM (Ho et al., 2020) when data entropy $H({\bm{x}}_{0})$ is excluded, thereby suggesting that optimizing DDPM’s variational loss directly tightens this explicit upper bound on the discrepancy between trued data and model-generated data.
@@
Noise model prediction error. The second source of errors stems from the noise prediction model. Because the model’s estimate ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ only approximates the true noise $\bm{\epsilon}$}, as guaranteed in principle by the Universal Approximation Theorem (Hornik et al., 1989; Nielsen, 2015; Yarotsky, 2022). Let $\Delta{\bm{\epsilon}}_{t}({\bm{x}}_{t})={\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ represent this prediction error. Based on Eq. (5), any discrepancy in ${\bm{\epsilon}}_{\theta}$ induces a corresponding deviation in the predicted mean at step $t$:
@@ A.3 The Discrepancy between DDPM and DDIM in the Forward and Reverse Process
As shown in Eq. (2), the forward process in DDPM (Ho et al., 2020) can be formulated as:
@@
For DDIM (Song et al., 2021), the forward process is identical to that in Eq. (A.16), while its deterministic reverse process (with $\eta=0$) is defined as follows:
@@ A.6 Examples of YOLO Predictions of Fingertips
Figure A3 presents representative examples of YOLO predictions on the RealHand dataset. The YOLO detector accurately identifies fingertips across diverse hand poses and orientations, including both dorsal and palmar views. However, it may still fail to recognize anatomically implausible configurations, such as hands with duplicated thumbs (e.g., the penultimate sub-figure of the first row and the last sub-figure of the last row in Fig. 3(a)) or hands with five middle fingers but without a thumb (e.g., the last sub-figure of the first row and the penultimate sub-figure of the last row in Fig. 3(a)). These abnormal samples are still detected as five-finger hands and therefore remain indistinguishable from anatomically valid ones in terms of finger counts.
Figure 3(b) shows several counting-hallucinated samples where YOLO predicts incorrect fingertip counts (e.g., 3, 4, 6, or 7 fingertips). Figure 3(c) further illustrates several rare edge cases (<1%) in which the detector fails to identify fingertips with unclear boundaries, cracked fingers, or floating artifacts. Such rare cases can be partly mitigated by tuning YOLO’s inference hyperparameters (e.g., IoU and confidence thresholds); however, doing so may compromise detection performance on the majority of normal samples. Since a consistent detection setup is applied across all evaluations on the RealHand dataset, the presence of these rare cases does not affect the validity of our overall conclusions, even though addressing them remains a challenging and low-reward effort.

### v2

@@ A.1 Poof of the Bound of Total Accumulated Distributional Error in Ancestral Sampling
Then, according to the Data Processing Inequality (cover2012elements), marginalizing (or discarding some random variables) will not increase the KL divergence:
@@
where equality (=) would strictly hold only in the ideal case where $q({\bm{x}}_{T})=p_{\theta}({\bm{x}}_{T})$ and $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})=q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})$ for all $t$ and ${\bm{x}}_{t}$, which means the diffusion model can accurately recover the underlying data distribution $q({\bm{x}}_{0})$. This upper bound for the overall KL divergence (${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{0})\|p_{\theta}({\bm{x}}_{0}))$) between true data and model-generated data is identical to the variational bound derived in DDPM (ddpm) when data entropy $H({\bm{x}}_{0})$ is excluded, thereby suggesting that optimizing DDPM’s variational loss directly tightens this explicit upper bound on the discrepancy between trued data and model-generated data.
@@
Noise model prediction error. The second source of errors stems from the noise prediction model. Because the model’s estimate ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ only approximates the true noise $\bm{\epsilon}$}, as guaranteed in principle by the Universal Approximation Theorem (hornik1989; nielsen2015neural; yarotsky2022universal). Let $\Delta{\bm{\epsilon}}_{t}({\bm{x}}_{t})={\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ represent this prediction error. Based on Eq. (5), any discrepancy in ${\bm{\epsilon}}_{\theta}$ induces a corresponding deviation in the predicted mean at step $t$:
@@ A.3 The Discrepancy between DDPM and DDIM in the Forward and Reverse Process
As shown in Eq. (2), the forward process in DDPM (ddpm) can be formulated as:
@@
For DDIM (ddim), the forward process is identical to that in Eq. (A.16), while its deterministic reverse process (with $\eta=0$) is defined as follows:
@@ A.6 Examples of YOLO Predictions of Fingertips
Figure A3 presents representative examples of YOLO predictions on the RealHand dataset. The YOLO detector accurately identifies fingertips across diverse hand poses and orientations, including both dorsal and palmar views. However, it may still fail to recognize anatomically implausible configurations, such as hands with duplicated thumbs (e.g., the penultimate sub-figure of the first row and the last sub-figure of the last row in Fig. 3(a)) or hands with five middle fingers but without a thumb (e.g., the last sub-figure of the first row and the penultimate sub-figure of the last row in Fig. 3(a)). These abcounting-valid samples are still detected as five-finger hands and therefore remain indistinguishable from anatomically valid ones in terms of finger counts.
Figure 3(b) shows several counting-hallucinated samples where YOLO predicts incorrect fingertip counts (e.g., 3, 4, 6, or 7 fingertips). Figure 3(c) further illustrates several rare edge cases (<1%) in which the detector fails to identify fingertips with unclear boundaries, cracked fingers, or floating artifacts. Such rare cases can be partly mitigated by tuning YOLO’s inference hyperparameters (e.g., IoU and confidence thresholds); however, doing so may compromise detection performance on the majority of counting-valid samples. Since a consistent detection setup is applied across all evaluations on the RealHand dataset, the presence of these rare cases does not affect the validity of our overall conclusions, even though addressing them remains a challenging and low-reward effort.
@@ A.7 Examples of Three Classes of Generated Hand Images
A.8 Extended Details of PTI-Hand experiments
In the PTI-Hand experiments, we adopt 10 different hand-generation prompts to prompt Stable-Diffusion-3.5-Medium. The prompts are show as follows.
  1. 1.
    A single human hand, front view, all fingers in frame and visible.
  2. 2.
    A complete human hand, seen from the front, showing all fingers clearly, minimal background.
  3. 3.
    An isolated single hand, viewed from directly above, showing all fingers, simple studio background.
  4. 4.
    Overhead shot of a person’s hand, entire hand in frame, all fingers visible, plain background.
  5. 5.
    One hand, top view, five fingers visible, simple background.
  6. 6.
    A single hand, front or back view, flat against a surface, all fingers in frame and fully visible.
  7. 7.
    A single human hand, top-down view, all fingers visible, resting on a plain white surface.
  8. 8.
    A single human hand seen from above, fingers spread naturally, minimal background.
  9. 9.
    A human hand, top view, palm facing downward, all fingers clearly in frame, neutral background.
  10. 10.
    A single human hand captured from overhead, fingers extended, plain background with soft lighting.
The definition of hallucinations assumes that the model generates patterns that were never observed during training. However, this assumption is difficult to verify for large-scale pretrained diffusion models, where the full training distribution is inaccessible. To make the quantification of counting hallucinations in such cases feasible, we adopt a conservative criterion: we treat any generated hand image containing six or more fingers as a counting-hallucinated sample. This assumption is well justified, while real images may depict fewer than five visible fingers due to occlusion, articulation, or viewpoint, it is highly unlikely for real data to contain hands with six or more fingers. Thus, finger counts $\geq$ 6 provide a reliable and empirically verifiable signal of counting hallucination. Generated examples can be found in Fig. A5.
A.9 Extended Details, Results, and Discussion of the Joint Diffusion Model
A.9.1 Implementation details in Table 3
In the JDM experiments reported in Table 3, we fine-tune the denoising UNet for 120k steps, compared to 80k steps for the vanilla LDM. This increased training budget is necessary because extending the latent space from three to six channels in JDM introduces a substantial mismatch from the pretrained latent distribution of LDM. As a result, additional optimization is required to close this distributional gap and allow JDM’s modeling capacity to fully materialize.
A.9.2 Fine-Tuning Pretrained Text-to-Image Models with Joint-Diffusion Training
Table A3: Comparison of failure rates and image quality between finetuning SD-1.5 and finetuning SD-1.5 with joint-diffusion training paradigm on RealHand. The first row (baseline) reports results from finetuning Stable-Diffusion-1.5 on RealHand data. The second row shows results obtained by freezing all intermediate layers of the baseline model and finetuning only the input and output layers with extended channel dimensions. Lower values indicate better performance.
Counting Hallucination Rate (CHR) $\downarrow$ Non-counting Failure Rate (NCFR) $\downarrow$ Fréchet Inception Distance (FID) $\downarrow$
Solver Name DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2
Sampling Steps 25 50 100 25 50 100 25 50 100 25 50 100 25 50 100 25 50 100
SD-1.5 0.3362 0.3404 0.3630 0.3396 0.3376 0.3362 0.0059 0.0032 0.0020 0.0050 0.0026 0.0018 61.51 57.71 57.70 58.53 57.57 56.69
SD-1.5 with joint-diffusion finetuning 0.2838 0.3222 0.4145 0.3053 0.3277 0.3558 0.0947 0.0511 0.0487 0.0549 0.0362 0.0327 70.65 79.22 102.39 70.80 72.84 83.07
Importantly, JDM should be viewed as a novel training paradigm for improving factual generation in diffusion models rather than as a superior inference strategy. On the one hand, JDM only extends the input and output channels of the UNet while keeping all sizes of intermediate layers unchanged, resulting in introduces a negligible inference-time overhead compared to vanilla LDM. On the other hand, extending a pretrained text-to-image model with additional mask channels disrupts the manifold geometry of the pretrained latent space and underlines the original denoising distribution, making direct finetuning such model nontrivial.
Despite this challenge, we still observe promising reductions in counting hallucinations when applying JDM to large-scale pretrained diffusion models under carefully controlled finetuning conditions. In practice, we find that jointly aligning the extended latent space in JDM with the highly rigid latent geometry of large pretrained models is challenging under the limited scale of RealHand. To strike a practical balance, we adopt a two-stage finetuning strategy: we first finetune a Stable-Diffusion-1.5 model on RealHand, then freeze all intermediate UNet layers and finetune only the input and output layers with extended channels. As shown in Table A3, even under this highly constrained finetuning setup, the joint-diffusion paradigm consistently yields lower counting hallucination rates at 25 and 50 sampling steps across both solvers.
Overall, the evidence presented in this work suggests that JDM is a promising direction for enforcing factual consistency in image generation. Lastly, we believe that exploring its behavior and performance under true large-scale training remains an important and compelling avenue for future research.
@@ A.9.2 Fine-Tuning Pretrained Text-to-Image Models with Joint-Diffusion Training
(a) Counting-valid samples (five valid fingers)
(b) Counting-hallucinated samples (more than 5 fingers)
(c) Non-counting failure samples (unsuitable for counting)
Figure A5: Examples of three classes of generated hand images using pretrained Stable-Diffusion-3.5-Medium (PTI-Hand). (a) Counting-valid: images with five clearly discernible fingers. (b) Counting-hallucinated: images with more than five valid fingers (e.g., 6 or 7). (c) Non-counting failures: images unsuitable for counting due to severe visual artifacts, such as distorted, cracked, partially visible, or floating fingers.